OpenStudy (anonymous):

What value of x makes the equation true? –4.9 = –5.7 + x A. x = –10.6 B. x = –0.8 C. x = 0.8 D. x = 10.6

OpenStudy (bishoppatton):

C.

OpenStudy (anonymous):

-4.9 = -5.7 + x add 5.7 to each side x = 0.8

OpenStudy (anonymous):

Solve for x: \[-4.9 = x - 5.7\] Reverse the equality in -4.9 = x - 5.7 in order to isolate x to the left hand side. \[x - 5.7 = -4.9\] Add +5.7 to both sides.. \[x = 0.8\] So your final answer is C. X = 0.8